OpenStudy (anonymous):

Solve for x. 5x - 6 = 4x + 8 x = -14 x = -2 x = 2 x = 14

OpenStudy (anonymous):

Solve for x. 5x - 6 = 4x + 8 x = -14 x = -2 x = 2 x = 14

OpenStudy (anonymous):

Arrange the x on one side and the numbers on the other and you should get your answer

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Move all the x terms to the left side and all the numbers to the right side

OpenStudy (anonymous):

Move the x terms to one side 5x-6=4x+8 5x-4x=8+6 x=14

OpenStudy (anonymous):

thank you